OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57

OpenStudy (anonymous):

If it makes it eayser do u want mutiple choice?

OpenStudy (anonymous):

Solve for x by getting rid of y 7x - 2y = 21 (4x + y = 57) x 2 (Mulitply second equation by 2) 7x - 2y = 21 8x + 2y = 114 (Add the two equations together getting rid of y) 15x = 135 x = 9